OpenStudy (anonymous):

If scores on an exam are normally distributed with a mean of 80% and a standard deviation of 10%, what would a score of 90% be in Z-scores (e.g., how many standard deviations is this score above or below the mean)? Z=-1.5 Z=-1.0 Z=0 Z=1.0 Z=1.5

OpenStudy (anonymous):

+1.0 right?

OpenStudy (amistre64):

z = (x-mean)/sd

OpenStudy (amistre64):

so yes, +1 sd looks fine

OpenStudy (anonymous):

ahahaha.. ti looks so easy to be real!! ;p get ready a quiz game begining ;)

OpenStudy (anonymous):

\[\bf z=\frac{ x-\mu }{ \sigma }\]Now plug in 90% for 'x', 80% for mean and 10% for S.D to get the z-score:\[\bf z = \frac{ 90-80 }{ 10 }=1\]Now look this up on a z-score table and so the probability of getting up to 90% would be:\[\bf P(z<1)=0.8413\]

OpenStudy (anonymous):

oh u just wanted to know the z-score..lol

OpenStudy (amistre64):

lol, ... the answer to a question that was never asked ;)

OpenStudy (anonymous):

genius you are too genius for these questions ;p... you are for next level questions just wait ;p

OpenStudy (anonymous):

@amistre64 I always assume that a z-score question is asking for the probability as well since it's usually the case with most questions...lol